OpenStudy (anonymous):

How would I factor 28y - 16

OpenStudy (anonymous):

\[28y-16=4(7y-4)\]

OpenStudy (anonymous):

That's what it would be factored?

OpenStudy (anonymous):

That's what it would be factored, but in order to get that, you have to find the greatest common factor of both 28 and 16. In this case, that happens to be 4.

OpenStudy (anonymous):

I understand it. Thank you for the help!

OpenStudy (anonymous):

Once you figure out that the greatest common factor of 28 and 16 is 4, then you write the 4 out front, start your set of parentheses, then ask yourself what 28y/4 would be. That answer is 7y, which is where that came from. Then bring down the - sign. Then ask yourself what 16/4 is, which is where the second 4 comes in. Then close your parentheses and you are done!

OpenStudy (anonymous):

Could you help me with one more? I don't get this one. X^2-9y^2

OpenStudy (anonymous):

I don't understand how to factor it

OpenStudy (anonymous):

Sure, are they two different variables? one is x and one is y?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok, if they have no common factors and no common variables, then it cannot be factored

OpenStudy (anonymous):

unless they want you to take a 1 out, in which case it would be 1(the original expression)

OpenStudy (anonymous):

So it can't be factored?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

That's not true. It can be factored.

OpenStudy (anonymous):

\[x^2-9y^2=x^2-(3y)^2=(x-3y)(x+3y)\]

OpenStudy (anonymous):

Sorry, I thought you were looking for a constant or single variable